Water is lifted out of a well 30 m deep by a motor of power 1.00 hp. Assuming 90% efficiency, how many gallons of water are lifted in 1 minute? [Hint: one gallon of pure water has a mass of 3.8 kg.
Answer #1

power is given by

P = \frac{mgh}{t}

0.90*1*746 = \frac{m*9.8*30}{60}

m = 137 kg

m = 36 Gallon